How many gallons of a fruit punch containing 20% alcohol should be mixed with 2 gallons of fruit punch containing 15% alcohol to obtain a fruit punch mixture that is 18% alcohol?

Answer by stanbon(75887)

How many gallons of a fruit punch containing 20% alcohol should be mixed with 2 gallons of fruit punch containing 15% alcohol to obtain a fruit punch mixture that is 18% alcohol?
------
Equation:
alc + alc = alc
0.20x + 0.15*2 = 0.18(x+2)
-----
20x + 15*2 = 18x + 18*2
2x = 3*2
x = 3 gallons (amt. of 20% solution needed)
===================
Cheers,
Stan H.
===================